# Amortized Variational Inference for Partial-Label Learning: A Probabilistic Approach to Label Disambiguation

### v1

@@ Abstract
Code and datahttps://github.com/mathefuchs/vi-pll
@@ 1 Introduction
Real-world datasets are prone to noise and labeling uncertainty. For instance, different annotators may assign conflicting class labels to the same instance. Such ambiguity frequently arises in applications like web mining (Guillaumin, Verbeek, and Schmid 2010; Zeng et al. 2013) and audio classification (Briggs, Fern, and Raich 2012). Although data cleaning can mitigate these issues, it is often time-consuming and resource-intensive. Partial-label learning (PLL; Zhang, Yu, and Tang 2017; Lv et al. 2020; Wang et al. 2022) provides a principled approach to learning from such data without requiring manual label cleaning. In PLL, each instance is associated with a set of candidate labels, exactly one of which is the unknown true label. PLL algorithms facilitate the training of multi-class classifiers in this weakly-supervised setting.
Recent state-of-the-art methods (Xu et al. 2021; Tian et al. 2024; Yang et al. 2025) leverage deep learning to optimize surrogate loss functions, such as the minimum loss (Lv et al. 2020), contrastive loss (Wang et al. 2022), or discriminative loss (Yang et al. 2025); and to refine candidate label sets through heuristic strategies, including importance reweighting (Feng et al. 2020), confidence thresholds (Xu et al. 2023), or label smoothing (Gong, Bisht, and Xu 2024). In contrast, earlier PLL methods (Jin and Ghahramani 2002; Liu and Dietterich 2012) approximate the posterior distribution over true labels directly—typically through expectation-maximization—rather than relying on heuristic refinements. However, these methods are computationally intensive and hardly scale.
@@ 3.1 Partial-Label Learning
PLL has received increasing attention over the past decades. Most existing approaches adapt standard supervised classification algorithms to the PLL context. Examples include nearest-neighbor methods (Hüllermeier and Beringer 2005; Zhang and Yu 2015; Fuchs, Kalinke, and Böhm 2025), support-vector machines (Nguyen and Caruana 2008; Cour, Sapp, and Taskar 2011; Yu and Zhang 2017), and label propagation strategies (Zhang and Yu 2015; Zhang, Zhou, and Liu 2016; Xu, Lv, and Geng 2019; Wang, Li, and Zhang 2019; Feng and An 2019).
Recent state-of-the-art methods (Tian et al. 2024; Yang et al. 2025; Fuchs and Kalinke 2025a) employ deep learning to train a multi-class classifier and iteratively refine the candidate sets. Since ground-truth labels are unavailable, optimization is performed using surrogate loss functions. For example, Lv et al. (2020); Feng et al. (2020) propose the minimum loss formulation, Xu et al. (2021) introduce a self-training strategy based on pseudo-labels, Zhang et al. (2022); Fuchs and Kalinke (2025b) leverage the magnitudes of class activation values, Wang et al. (2022) use ideas from contrastive learning, Xu et al. (2023) utilize level sets to iteratively remove incorrect labels from the candidate sets, Tian et al. (2024) propose a cross-model selection strategy, where multiple models are trained simultaneously, and Yang et al. (2025) introduce a pseudo-labeling framework based on the feature representations. While these methods use heuristics and surrogate losses to refine the candidate sets, our method directly approximates the true label posterior, offering a more principled solution to uncovering the hidden ground truth.
We note that several existing methods incorporate probabilistic components, such as modeling labels with a Dirichlet distribution (Xu et al. 2021; Fuchs and Kalinke 2025b). However, unlike our approach, these methods employ the Dirichlet model as an auxiliary mechanism, either as a regularization term or as an enhancement to classifier training, rather than as the core of their method.
Our approach is closest to the expectation-maximization strategies by Jin and Ghahramani (2002) and Liu and Dietterich (2012). However, these methods are computationally expensive, which limits their application on real-world datasets. In contrast, our method leverages NNs to amortize inference by directly predicting variational parameters from input data, enabling efficient and scalable learning.
@@ 3.2 Variational Inference
Variational methods offer a principled framework for approximate Bayesian inference by formulating posterior estimation as an optimization problem. Early methods (Jordan et al. 1999; Attias 1999; Beal 2003) introduce mean-field approximations and EM algorithms for latent variable models. These approaches typically rely on model-specific derivations and coordinate ascent updates. Kingma and Welling (2014) introduce variational auto-encoders (VAE) and amortized VI, which employ a NN (the encoder) to predict the variational parameters directly from input data. They also make use of the reparameterization trick to enable backpropagation through stochastic variables, facilitating scalable and efficient inference. Rezende, Mohamed, and Wierstra (2014) propose a similar approach in the context of deep latent Gaussian models. Subsequent extensions include VI with normalizing flows (Rezende and Mohamed 2015) and a model-agnostic optimization formulation (Ranganath, Gerrish, and Blei 2014). Sohn, Lee, and Yan (2015) introduce conditional variational auto-encoders (CVAE), which extend VAEs by conditioning the variational parameters on additional input data. Our method builds on this formulation and is detailed in the next section. An overview of recent developments in amortized VI in general is that of Margossian and Blei (2024).
@@ 4.1 Optimization Objective
where $f_{\phi}:\operatorname{\mathcal{X}}\times 2^{\operatorname{\mathcal{Y}}}\to\mathbb{R}^{k}_{\geq 0}$ denotes a NN that maps input features and candidate label sets to non-negative parameter vectors, and $k$ is the number of classes. Non-negativity is enforced by applying a softplus activation function (Glorot, Bordes, and Bengio 2011) in the output layer of the network, while the Dirichlet distribution enables modeling uncertainty over class label assignments (Jøsang 2016; Sensoy, Kaplan, and Kandemir 2018). Since $\alpha_{\phi}\geq 1$, $\mathrm{Dir}(y;\alpha_{\phi})$ only has a single mode, which eases its optimization.
@@
We argue that (4) is beneficial in our setting as it explicitly allows modeling prior information $P(Y)$ on the class labels as well as a generative model of the observed features $P(X\mid Y)$ given label information. Additionally, in (4), the unobserved variable $Y$ is not dependent on any other variables, which eases its modeling. In contrast, existing work (Liu and Dietterich 2012; Feng et al. 2020) relies on the factorization in (3). Our experiments in Section 5.3 confirm our argument in favor of (4).
@@
Using (4), we have $P_{\theta,\gamma}(Y\mid X,S)=P(Y)P_{\theta,\gamma}(X\mid Y)P(S\mid X,Y)/P(X,S)$, where in Section 4.2, we model $P_{\theta,\gamma}(X\mid Y)$ with a conditional variational auto-encoder (CVAE; Kingma and Welling 2014; Sohn, Lee, and Yan 2015), Section 4.3 elaborates on the prior term $p(y)$, and Section 4.4 details the candidate set distribution $p(s\mid x,y)$.
@@
where $\arg\min_{\phi,\theta,\gamma}\operatorname{\mathcal{L}}(\phi,\theta,\gamma)=\arg\max_{\phi,\theta,\gamma}\text{$\beta$-ELBO}$, for $\beta=1$. The scaling parameter $\beta\in(0,1]$ allows weighting (10) for more flexibility (Higgins et al. 2017).
In practice, we replace $\operatorname{\mathbb{E}}_{(x,s)\sim P_{XS}}$ by a sample average over mini-batches $(x_{i},s_{i})\in\operatorname{\mathcal{D}}_{\mathrm{mb}}\subseteq\operatorname{\mathcal{D}}$. Similarly, we replace $\operatorname{\mathbb{E}}_{y\sim q_{\phi}(y\mid x,s)}$ by taking a sample average of $b$ Monte Carlo samples $y_{i}\sim q_{\phi}(y\mid x,s)$ for $i\in[b]$. Sampling from the variational distribution is straightforward, since $q_{\phi}(y\mid x,s)=\mathrm{Dir}(y;\alpha_{\phi})$ and $p(y)=\mathrm{Dir}(y;\alpha^{\pi})$ (compare Section 4.3) admit a closed-form expression for their KL divergence in (10) (Penny 2001).
@@
Factorizing the joint density $P(X,Y,S)$ via either (3) or (4) has the interpretation of a directed acyclic graph (DAG). Specifically, the factorization in (3) corresponds to the DAG in Figure 1(a) (which is the standard in the literature; see Cour, Sapp, and Taskar 2011) and the factorization in (4) to that in Figure 1(b), which we use in our work. Proposition 4.3 elaborates on their equivalence, that is, having offline data only, one cannot distinguish the two. Both models explain the observed data equally well.
@@
Recall from (Verma and Pearl 1990, Theorem 1) that two DAGs are Markov equivalent if and only if they have (1) the same skeleton, that is, the same edges ignoring direction, and (2) the same v-structures, that is, the same set of nodes $A\rightarrow C\leftarrow B$, where $A$ and $B$ are not connected. In our setting, (1) and (2) are satisfied. ∎
@@
Figure 1: Different DAGs used in PLL. Figure 1(a) shows the standard model from the literature (Cour, Sapp, and Taskar 2011). Figure 1(b) shows the Markov-equivalent model (compare Proposition 4.3) that we use in our work.
@@ 4.2 Generative Model $p_{\theta,\gamma}(x\mid y)$
where we use the common log-sum-exp trick for numerical stability (Blanchard, Higham, and Higham 2021). After warming-up the CVAE, we set $\sigma$ in (19) to an exponential moving average of the observed RMSE reconstruction loss.
@@ 4.4 Candidate Set Distribution $p(s\mid x,y)$
where, in $(i)$, we assume that $x$ and $s$ are conditionally independent given $y$, which is a common assumption in the literature (Liu and Dietterich 2012; Feng et al. 2020). In $(ii)$, we express $p(s\mid y)$ as the amount of labeling information that agrees with the candidate set $s$. Therefore, (23) acts as a regularization term enforcing that the constraints from the candidate sets are satisfied. Proposition 4.4 demonstrates that this is a valid mass function.
@@ 4.5 Proposed Algorithm
Phase 2 (Lines 4–10) is the warm-up phase for the CVAE in which we minimize the reconstruction and regularization losses in (20) and (18), respectively, using the labeling vectors $\tilde{y}_{i}$. We use mini-batches and train for $T_{\mathrm{w}}=500$ epochs using the Adam optimizer (Kingma and Ba 2015).
@@ 5 Experiments
Section 5.1 summarizes all PLL methods that we compare against, Section 5.2 describes our experimental setup including the datasets and candidate generation strategies used, and Section 5.3 shows our main findings.
@@ 5.1 Competitors
Besides our method ViPll and a variant of it containing ablations, we consider nine established benchmarks from the literature. These are PlKnn (Hüllermeier and Beringer 2005) and PlEcoc (Zhang, Yu, and Tang 2017), as well as the state-of-the-art deep learning methods Proden (Lv et al. 2020), Valen (Xu et al. 2021), Cavl (Zhang et al. 2022), PiCO (Wang et al. 2022), Pop (Xu et al. 2023), CroSel (Tian et al. 2024), and Cel (Yang et al. 2025). For a fair comparison, we use the same base models for all approaches, that is, an MLP with ReLU activation and batch normalization. For the colored image datasets, we use the pre-trained Blip2 model (Li et al. 2023) to extract 768-dimensional feature vectors.
@@ 5.2 Experimental Setup
As is common in the PLL literature (Lv et al. 2020; Xu et al. 2023), we use real-world PLL datasets as well as supervised multi-class classification datasets with added candidates. Table 1 summarizes the dataset characteristics. The real-world PLL datasets include the bird-song (Briggs, Fern, and Raich 2012), lost (Cour, Sapp, and Taskar 2011), mir-flickr (Huiskes and Lew 2008), msrc-v2 (Liu and Dietterich 2012), and yahoo-news dataset (Guillaumin, Verbeek, and Schmid 2010). The supervised multi-class classification datasets include the mnist (LeCun, Cortes, and Burges 1999), fmnist (Xiao, Rasul, and Vollgraf 2018), kmnist (Clanuwat et al. 2018), cifar10 (Krizhevsky 2009), and cifar100 dataset (Krizhevsky 2009).
@@
To obtain partial labels for the five supervised datasets, we add candidate labels using the common instance-dependent generation strategy (Xu et al. 2021; Yang et al. 2025) as follows. One first trains a supervised MLP classifier $g:\operatorname{\mathcal{X}}\to\operatorname{\Delta^{k}}$. Then, given an instance $x\in\operatorname{\mathcal{X}}$ with correct label $y\in\operatorname{\mathcal{Y}}$, a binomial flipping probability of $\xi_{1}(x,\bar{y})=g_{\bar{y}}(x)/\max_{y^{\prime}\in\operatorname{\mathcal{Y}}\setminus\{\bar{y}\}}g_{y^{\prime}}(x)$ decides whether to add the incorrect label $\bar{y}\neq y$ to the candidate set $s$, that is, one samples $w_{\bar{y}}\sim\mathcal{U}(0,1)$ for each $\bar{y}\neq y$ and adds it to $x$’s candidate set $s$ if $w_{\bar{y}}\leq\xi_{1}(x,\bar{y})$.

### v2

@@ Abstract
Partial-Label Learning, Weakly Supervised Learning, Amortized Variational Inference
@@ 1 Introduction
Real-world datasets are prone to noise and labeling uncertainty. For instance, different annotators may assign conflicting class labels to the same instance. Such ambiguity frequently arises in applications like web mining (Guillaumin et al., 2010; Zeng et al., 2013) and audio classification (Briggs et al., 2012). Although data cleaning can mitigate these issues, it is often time-consuming and resource-intensive. Partial-label learning (PLL; Zhang et al. 2017; Lv et al. 2020; Wang et al. 2022) provides a principled approach to learning from such data without requiring manual label cleaning. In PLL, each instance is associated with a set of candidate labels, exactly one of which is the unknown true label. PLL algorithms facilitate the training of multi-class classifiers in this weakly-supervised setting.
Recent state-of-the-art methods (Xu et al., 2021; Tian et al., 2024; Yang et al., 2025) leverage deep learning to optimize surrogate loss functions, such as the minimum loss (Lv et al., 2020), contrastive loss (Wang et al., 2022), or discriminative loss (Yang et al., 2025); and to refine candidate label sets through heuristic strategies, including importance reweighting (Feng et al., 2020), confidence thresholds (Xu et al., 2023), or label smoothing (Gong et al., 2024). In contrast, earlier PLL methods (Jin & Ghahramani, 2002; Liu & Dietterich, 2012) approximate the posterior distribution over true labels directly—typically through expectation-maximization—rather than relying on heuristic refinements. However, these methods are computationally intensive and hardly scale.
@@ 3.1 Partial-Label Learning
PLL has received increasing attention over the past decades. Most existing approaches adapt standard supervised classification algorithms to the PLL context. Examples include nearest-neighbor methods (Hüllermeier & Beringer, 2005; Zhang & Yu, 2015; Fuchs et al., 2025), support-vector machines (Nguyen & Caruana, 2008; Cour et al., 2011; Yu & Zhang, 2017), and label propagation strategies (Zhang & Yu, 2015; Zhang et al., 2016; Xu et al., 2019; Wang et al., 2019; Feng & An, 2019).
Recent state-of-the-art methods (Tian et al., 2024; Yang et al., 2025; Fuchs & Kalinke, 2025a) employ deep learning to train a multi-class classifier and iteratively refine the candidate sets. Since ground-truth labels are unavailable, optimization is performed using surrogate loss functions. For example, Lv et al. (2020); Feng et al. (2020) propose the minimum loss formulation, Xu et al. (2021) introduce a self-training strategy based on pseudo-labels, Zhang et al. (2022); Fuchs & Kalinke (2025b) leverage the magnitudes of class activation values, Wang et al. (2022) use ideas from contrastive learning, Xu et al. (2023) utilize level sets to iteratively remove incorrect labels from the candidate sets, Tian et al. (2024) propose a cross-model selection strategy, where multiple models are trained simultaneously, and Yang et al. (2025) introduce a pseudo-labeling framework based on the feature representations. While these methods use heuristics and surrogate losses to refine the candidate sets, our method directly approximates the true label posterior, offering a more principled solution to uncovering the hidden ground truth.
We note that several existing methods incorporate probabilistic components, such as modeling labels with a Dirichlet distribution (Xu et al., 2021; Fuchs & Kalinke, 2025b). However, unlike our approach, these methods employ the Dirichlet model as an auxiliary mechanism, either as a regularization term or as an enhancement to classifier training, rather than as the core of their method.
Our approach is closest to the expectation-maximization strategies by Jin & Ghahramani (2002) and Liu & Dietterich (2012). However, these methods are computationally expensive, which limits their application on real-world datasets. In contrast, our method leverages NNs to amortize inference by directly predicting variational parameters from input data, enabling efficient and scalable learning.
@@ 3.2 Variational Inference
Variational methods offer a principled framework for approximate Bayesian inference by formulating posterior estimation as an optimization problem. Early methods (Jordan et al., 1999; Attias, 1999; Beal, 2003) introduce mean-field approximations and EM algorithms for latent variable models. These approaches typically rely on model-specific derivations and coordinate ascent updates. Kingma & Welling (2014) introduce variational auto-encoders (VAE) and amortized VI, which employ a NN (the encoder) to predict the variational parameters directly from input data. They also make use of the reparameterization trick to enable backpropagation through stochastic variables, facilitating scalable and efficient inference. Rezende et al. (2014) propose a similar approach in the context of deep latent Gaussian models. Subsequent extensions include VI with normalizing flows (Rezende & Mohamed, 2015) and a model-agnostic optimization formulation (Ranganath et al., 2014). Sohn et al. (2015) introduce conditional variational auto-encoders (CVAE), which extend VAEs by conditioning the variational parameters on additional input data. Our method builds on this formulation and is detailed in the next section. An overview of recent developments in amortized VI in general is that of Margossian & Blei (2024).
@@ 4.1 Optimization Objective
where $f_{\phi}:\operatorname{\mathcal{X}}\times 2^{\operatorname{\mathcal{Y}}}\to\mathbb{R}^{k}_{\geq 0}$ denotes a NN that maps input features and candidate label sets to non-negative parameter vectors, and $k$ is the number of classes. Non-negativity is enforced by applying a softplus activation function (Glorot et al., 2011) in the output layer of the network, while the Dirichlet distribution enables modeling uncertainty over class label assignments (Jøsang, 2016; Sensoy et al., 2018). Since $\alpha_{\phi}\geq 1$, $\mathrm{Dir}(y;\alpha_{\phi})$ only has a single mode, which eases its optimization.
@@
We argue that (4) is beneficial in our setting as it explicitly allows modeling prior information $P(Y)$ on the class labels as well as a generative model of the observed features $P(X\mid Y)$ given label information. Additionally, in (4), the unobserved variable $Y$ is not dependent on any other variables, which eases its modeling. In contrast, existing work (Liu & Dietterich, 2012; Feng et al., 2020) relies on the factorization in (3). Our experiments in Section 5.3 confirm our argument in favor of (4).
@@
Using (4), we have $P_{\theta,\gamma}(Y\mid X,S)=P(Y)P_{\theta,\gamma}(X\mid Y)P(S\mid X,Y)/P(X,S)$, where in Section 4.2, we model $P_{\theta,\gamma}(X\mid Y)$ with a conditional variational auto-encoder (CVAE; Kingma & Welling 2014; Sohn et al. 2015), Section 4.3 elaborates on the prior term $p(y)$, and Section 4.4 details the candidate set distribution $p(s\mid x,y)$.
@@
where $\arg\min_{\phi,\theta,\gamma}\operatorname{\mathcal{L}}(\phi,\theta,\gamma)=\arg\max_{\phi,\theta,\gamma}\text{$\beta$-ELBO}$, for $\beta=1$. The scaling parameter $\beta\in(0,1]$ allows weighting (10) for more flexibility (Higgins et al., 2017).
In practice, we replace $\operatorname{\mathbb{E}}_{(x,s)\sim P_{XS}}$ by a sample average over mini-batches $(x_{i},s_{i})\in\operatorname{\mathcal{D}}_{\mathrm{mb}}\subseteq\operatorname{\mathcal{D}}$. Similarly, we replace $\operatorname{\mathbb{E}}_{y\sim q_{\phi}(y\mid x,s)}$ by taking a sample average of $b$ Monte Carlo samples $y_{i}\sim q_{\phi}(y\mid x,s)$ for $i\in[b]$. Sampling from the variational distribution is straightforward, since $q_{\phi}(y\mid x,s)=\mathrm{Dir}(y;\alpha_{\phi})$ and $p(y)=\mathrm{Dir}(y;\alpha^{\pi})$ (compare Section 4.3) admit a closed-form expression for their KL divergence in (10) (Penny, 2001).
@@
Factorizing the joint density $P(X,Y,S)$ via either (3) or (4) has the interpretation of a directed acyclic graph (DAG). Specifically, the factorization in (3) corresponds to the DAG in Figure 1(a) (which is the standard in the literature; see Cour et al., 2011) and the factorization in (4) to that in Figure 1(b), which we use in our work. Proposition 4.3 elaborates on their equivalence, that is, having offline data only, one cannot distinguish the two. Both models explain the observed data equally well.
@@
Recall from (Verma & Pearl, 1990, Theorem 1) that two DAGs are Markov equivalent if and only if they have (1) the same skeleton, that is, the same edges ignoring direction, and (2) the same v-structures, that is, the same set of nodes $A\rightarrow C\leftarrow B$, where $A$ and $B$ are not connected. In our setting, (1) and (2) are satisfied. ∎
@@
Figure 1: Different DAGs used in PLL. Figure 1(a) shows the standard model from the literature (Cour et al., 2011). Figure 1(b) shows the Markov-equivalent model (compare Proposition 4.3) that we use in our work.
@@ 4.2 Generative Model $p_{\theta,\gamma}(x\mid y)$
where we use the common log-sum-exp trick for numerical stability (Blanchard et al., 2021). After warming-up the CVAE, we set $\sigma$ in (19) to an exponential moving average of the observed RMSE reconstruction loss.
@@ 4.4 Candidate Set Distribution $p(s\mid x,y)$
where, in $(i)$, we assume that $x$ and $s$ are conditionally independent given $y$, which is a common assumption in the literature (Liu & Dietterich, 2012; Feng et al., 2020). In $(ii)$, we express $p(s\mid y)$ as the amount of labeling information that agrees with the candidate set $s$. Therefore, (23) acts as a regularization term enforcing that the constraints from the candidate sets are satisfied. Proposition 4.4 demonstrates that this is a valid mass function.
@@ 4.5 Proposed Algorithm
Phase 2 (Lines 4–10) is the warm-up phase for the CVAE in which we minimize the reconstruction and regularization losses in (20) and (18), respectively, using the labeling vectors $\tilde{y}_{i}$. We use mini-batches and train for $T_{\mathrm{w}}=500$ epochs using the Adam optimizer (Kingma & Ba, 2015).
@@ 5 Experiments
Section 5.1 summarizes all PLL methods that we compare against, Section 5.2 describes our experimental setup including the datasets and candidate generation strategies used, and Section 5.3 shows our main findings. Our code, data, and candidate generation strategy is openly available.1
@@ 5.1 Competitors
Besides our method ViPll and a variant of it containing ablations, we consider nine established competitors. These are PlKnn (Hüllermeier & Beringer, 2005) and PlEcoc (Zhang et al., 2017), as well as the state-of-the-art deep learning methods Proden (Lv et al., 2020), Valen (Xu et al., 2021), Cavl (Zhang et al., 2022), PiCO (Wang et al., 2022), Pop (Xu et al., 2023), CroSel (Tian et al., 2024), and Cel (Yang et al., 2025). For a fair comparison, we use the same base models for all approaches, that is, an MLP with ReLU activation and batch normalization. For the colored image datasets, we use the pre-trained Blip2 model (Li et al., 2023) to extract 768-dimensional feature vectors.
@@ 5.2 Experimental Setup
As is common in the PLL literature (Lv et al., 2020; Xu et al., 2023), we use real-world PLL datasets as well as supervised multi-class classification datasets with added candidates. Table 1 summarizes the dataset characteristics. The real-world PLL datasets include the bird-song (Briggs et al., 2012), lost (Cour et al., 2011), mir-flickr (Huiskes & Lew, 2008), msrc-v2 (Liu & Dietterich, 2012), and yahoo-news dataset (Guillaumin et al., 2010). The supervised multi-class classification datasets include the mnist (LeCun et al., 1999), fmnist (Xiao et al., 2018), kmnist (Clanuwat et al., 2018), cifar10 (Krizhevsky, 2009), and cifar100 dataset (Krizhevsky, 2009).
@@
To obtain partial labels for the five supervised datasets, we add candidate labels using the common instance-dependent generation strategy (Xu et al., 2021; Yang et al., 2025) as follows. One first trains a supervised MLP classifier $g:\operatorname{\mathcal{X}}\to\operatorname{\Delta^{k}}$. Then, given an instance $x\in\operatorname{\mathcal{X}}$ with correct label $y\in\operatorname{\mathcal{Y}}$, a binomial flipping probability of $\xi_{1}(x,\bar{y})=g_{\bar{y}}(x)/\max_{y^{\prime}\in\operatorname{\mathcal{Y}}\setminus\{\bar{y}\}}g_{y^{\prime}}(x)$ decides whether to add the incorrect label $\bar{y}\neq y$ to the candidate set $s$, that is, one samples $w_{\bar{y}}\sim\mathcal{U}(0,1)$ for each $\bar{y}\neq y$ and adds it to $x$’s candidate set $s$ if $w_{\bar{y}}\leq\xi_{1}(x,\bar{y})$.